light of frequency 5.0 * 14 h e r t z library electrons with energy .31 x 10 to the power - 19 joules from a certain metallic surface . Calculate the wavelength of ultraviolet light which liberates electrons of energy 8.93 10* 19 J from thesame surface
1
Expert's answer
2020-03-16T13:15:39-0400
"hf=A+E"

"hf'=A+E'"

"h(f'-f)=E'-E"

"(f'-5\cdot10^{14})=\frac{8.93\cdot10^{-19}-0.31\cdot10^{-19}}{6.63\cdot10^{-34}}"

"f'=1.3\cdot10^{15}Hz"